Solve the following systems of equations and state if any lines are parallel

1. y + 4x + 5 = 2
y + 3x = 2
2. y + 7x + 6 = 0
y = 2x + 30
3. -6x + 0 = y
6y -12x + 126 = 30
4. 9x - 8 = y
8x - 4 = y
5. y - 8x = 7
y - 8x - 7 = 0
6. y - 3x - 3 = 0
y - 5x + 15 = 0
1. y - 5x + 0 = 0
y + 1x = -24
2. y = 1x + 5
-7x + 45 = y
3. y - x -1 = 0
y = -6x - 41
4. y - 3x - 18 = -10
-6x + 71 = y
5. 0x - 3 = y
y + 8x + 27 = 0
6. 5x + 5y = y
y = -4x + 50

Answer by math-vortex(648)

Hi, there--
.
Here is a solution to #5
.
0x - 3 = y ---------------------> y = -3
y + 8x + 27 = 0 ----------------> 7 + 8x +27 = 0
.
Use the substitution method.
.
Substitute -3 for y in the second equation.
y + 8x + 27 = 0
(-3) + 8x + 27 = 0
.
Simplify and solve for x.
24 + 8x = 0
.
Subtract 24 from both sides of the equation.
8x = -24
.
Divide both sides of the equation by 8.
x = -3
.
The solution to this system is the ordered pair (-3,-3).
.
Please post one problem at a time.
.
Thank you,
.
Ms.Figgy
math.in.the.vortex@gmail.com